OpenStudy (anonymous):

Jacob has 3090 jelly beans. He gave 1295 to his teacher and he ate 277. How many of the jelly beans are blue?

OpenStudy (anonymous):

Well, being as the question states nothing of any color, I am going to assume that they are ALL blue. And if they are ALL blue, then he must have 1518 jelly beans left, all of which must be blue.

OpenStudy (anonymous):

Great, thank you!

OpenStudy (gorv):

you ate = 277 given=1295 so total spent=277+1295=1572

OpenStudy (gorv):

left with =3090-1572

OpenStudy (anonymous):

@Lparr You are VERY welcome!